Circular No. 6620

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 J. L. Tonry, E. M. Hu, and L. L. Cowie, Institute for Astronomy,

University of Hawaii; and R. G. McMahon, Institute of Astronomy, Cambridge, report: “Spectra (range 375–930 nm; resolution 1.0 nm FWHM) of the optical transient (IAUC 6584) associated with GRB 970228 (IAUC 6572) were obtained on the Keck II 10-m telescope (+ LRIS) in six exposures on Mar. 31.25 (500 and 1000 s), Apr. 1.25 (1000 and 1000 s), and Apr. 2.25 UT (1000 and 1000 s) totaling 5500 s under photometric conditions and ∼ 0′′.8-FWHM seeing. The slit (1′′ × 2′.9 at p.a. 86^{o}.4) was centered on a nearby star (cf. IAUC 6588) 2′′.9 east of the transient. The transient (α = 5^{h}01^{m}46^{s}.65, δ = +11^{o}46′53′′.6, equinox 2000.0; estimated from the HST WFPC2 image, IAUC 6606) is weakly detected in the spectrogram with no ‘obvious’ emission lines. The nearby star, at position end figures 46^{s}.46, 53′′.6, has a mid-K-type spectrum and no evidence for the TiO bands reported on IAUC 6588. A faint emission-line galaxy with redshift 0.638 ± 0.005 is detected 7′′.3 east of the transient (position end figures 47^{s}.15, 54′′.5). A second emission-line galaxy (position end figures 42^{s}.28, 50′′.7) has a redshift of 0.393 ± 0.005. The data are available at

http://www.ifa.hawaii/faculty/hu/grb.html or by anonymous ftp to
hubble.ifa.hawaii.edu:/pub/grb970228.”

COMET C/1995 O1 (HALE-BOPP) L. Jorda, Max-Planck-Institut für Aeronomie, Lindau; F. Colas, Bu-

reau des Longitudes; J. Lecacheux, Observatoire de Paris, Meudon; E. Frappa, Planetarium de St.-Etienne; A. Klotz, CESR, Toulouse; and S. Champenois, Laboratoire d’Astronomie de Lille, report: “The apparent rotation of the nucleus was clockwise on images obtained before Feb. 1997 (Kidger et al. 1996, Ap.J. 461, L119) and became distinctly anticlockwise during March, showing that the north pole is now directed toward the earth. The spiral jets appeared straight on visible and infrared images taken at Pic du Midi Observatory on Mar. 1 ± 2 days, showing that the line of sight crossed the equatorial plane of the nucleus at that date. This implies that the source of the bright jet observed since January is located in the northern hemisphere, contrary to what we announced in a preliminary analysis (IAUC 6560). This jet, observed in the hemisphere directed toward the sun, now has the appearance of concentric arcs with increasing curvature. A secondary dust jet is observed at p.a. 195^{o}–245^{o}.”

1997 April 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green